OpenStudy (calculusxy):

It is given that a + b = 6 and ab = 7. (a) Find the value of a^2 + b^2. Can you do this without finding values for a and b?

OpenStudy (calculusxy):

@jim_thompson5910

satellite73 (satellite73):

why not?

satellite73 (satellite73):

what is \((a+b)^2\)

OpenStudy (calculusxy):

a^2 + 2ab + b^2

satellite73 (satellite73):

you know all you need

OpenStudy (calculusxy):

So (6)^2 = 36 and 2(7) = 14 because (a + b) = 6 and we have (a + b)^2 and ab = 7 and 2ab^2

OpenStudy (calculusxy):

There's a second part to this question where I need to make a geometry word problem which would have to correspond to the question in part (a).

OpenStudy (calculusxy):

Do you have any suggestions for what I should do?

satellite73 (satellite73):

probably draw a square with side \(a+b\) i guess

satellite73 (satellite73):



OpenStudy (mathmale):

It's not clear whether or not you feel you have done the first part of this problem. I'd suggest you experiment. Note that a^2 + b^2 is not a perfect square, but you can make it into an equivalent expression that containts a perfect square. Suppose you begin with a^2 + b^2 and add 2ab to both sides: a^2 +2ab + b^2 -2ab is the same, algebraically, as a^2 + b^2. a^2 + 2ab + b^2 - 2ab = (a+b)^2 - 2ab. But you know that a+b=6 and ab=7. Substitute these knowns into (a+b)^2 - 2ab .

satellite73 (satellite73):

is this some time of math ed problem?

OpenStudy (calculusxy):

@satellite73 I am not sure what I am supposed to do w/ that square

satellite73 (satellite73):

fill in the areas

satellite73 (satellite73):

i meant that is my guess, not sure what else to do

satellite73 (satellite73):



OpenStudy (calculusxy):

Thank you so much! Can you help me with a couple of other problems?

satellite73 (satellite73):

i can try...